OpenStudy (anonymous):

Is the net force on a 3-kg free-falling object 10N?

OpenStudy (yash2651995):

^WHAT??

OpenStudy (yash2651995):

if there is no net force how will it still be in acceleration? net force is m.g for free falling body too.

OpenStudy (yash2651995):

3*9.8 N or roughly 3*10 N = 30 N

OpenStudy (unklerhaukus):

\[F=ma\]\[\quad=mg\]\[\qquad=3.00[\text{kg}]\cdot9.81[\text m / \text s^2]\]\[\qquad=3.00\times9.81[\text{kg}]\cdot[\text m / \text s^2]\]\[\qquad=29.4[\text{N}]\]